a. If Canace Company, with a break-even point at $520,000 of sales, has actual sales of $800,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number. 1. $ 2. % b. If the margin of safety for Canace Company was 45%, fixed costs were $1,947,825, and variable costs were 55% of sales, what was the amount of actual sales (dollars)? (Hint: Determine the break-even in sales dollars first.) $
Answer #1
margin of safety = total sale-breakeven point
margin of safety in dollars = 800000-520000 = 280000
margin of safety in % = margin of safety/sale
margin of safety in % = 280000/800000 = 35%
contribution margin ratio = 1-variable cost percentage
contribution margin ratio = 1-0.55 =0.45 (i.e. 45%)
Breakeven in dollars = fixed cost/contribution margin ratio
breakeven in dollars = 1947825/0.45 = 4328500
actual sale = 4328500/0.55 = 7870000

Please give a thumbs up if it is helpful & let me know if any doubt